Name: Q1. A car travels with a constant speed on a flat curve. The curve is circulating witha radius of 100 m and the coefficient of friction between the wheels and the surface of the road is 0.4. What is the maximum speed of the car so that it will stay on the road?
Answer #1

Using force balance on the car:

Fc = Ff

Centripetal force on car on the curve = friction force on car

m*ac = uk*N

m*Vmax^2/R = uk*m*g

Vmax = sqrt (uk*g*R)

Using given values:

Vmax = sqrt (0.4*9.8*100)

Vmax = 19.8 m/sec = max speed of car to stay on the road

Please Upvote.